How do you factor 2x3-10x2+40=8x
Answer by stanbon(75887)

factor 2x^3-10x^2+40=8x
----
Rearrange:
2x^3-10x^2-8x+40 = 0
---
2x^3-8x -10x^2+40 = 0
---
2x(x^2-4) -10(x^2-4) = 0
---
(x^2-4)(2x-10) = 0
x^2-4 = 0 or 2x-10 = 0
x = +-2 or x = 5
=====================
Cheers,
Stan H.
==============